Shift Summary, Eve 7/27/99

Peter Shanahan (shanahan@fnal.gov)
Wed, 28 Jul 1999 00:03:25 -0500

Shift Summary for Eve shift, 27/Jul/99

On Shift: PhG, PSh

Hours of beam available at all intensities: 7.8/8.5
Hours of beam available at nominal intensity: 7.5/8.5
Hours of E832 data (special runs not included): 7.3
Hours of special runs (laser scans, muon runs...): 0.1 (MUON)

Accounting of lost time: TeV startup, TeV fridge problems, a handful
of spills with ktevana errors.

========================================================================
Runs ended this shift:
Run 14043- 500k events - BEAM_MUON

Tapes brought to FCC this shift: NONE
(including tapes for online split)

========================================================================
Accesses during shift: NONE

========================================================================
Overview of shift:
E.g., Startup after 1.5 day shutdown. Muon run, then long BEAM832 run.

Phillipe labeled 90 tapes!

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14043 30 15:34 End muon run
15:45 Request hadron mode, ramp NM4AN to POS 1600
15:45 Fresh start for E832 mode
15:50 Just as we are ready to start, beam goes away due to
TeV corrector problems

14044 1 16:05 Beam returns. Request beamstop out, start run
~11 16:20 Complaints from SDAQ about RAW_L1_B1/SEM being ~20%
too low. No other rates back this up, except RAW_L1_B4,
REG_TOP. Problem goes away after a few spills.
~60 17:45 MU2 gives ~20K unknown crate errors. Fixed by kfrend.
18:50 DN installs new Ar-Ethane and CO2 bottles.
135 19:03 Change variable filter wheel setting to increase
quiescent laser light level.
170 19:54 Missing FFFFs in AZLAT crate --> kfrend.
202 SDAQ DCOR process claims DC4 oscillations. None seen.
OK on next cycle.
220 21:00 Called MCR to complain about 5-10% low intensity all
shift.
21:25 Mount QKA004 on ktev3 per onsplit request. Jim yells
at us anyway.
~270 22:06 Beam goes away due to TeV D1 fridge problem
284 22:24 Beam returns
348 23:52 Missing FFFFs in AZLAT crate --> kfrend.

========================================================================
Follow up Notes on Previous Shift

Drinking water was ordered by day shift.

========================================================================
Items to Follow up on Next Shift

Keep going with E832 running, skip NM4AN reversal.